# Are Lithium-Rich Giants Binaries?: A Radial Velocity Variability Analysis of 1,400 Giants

### v2

@@ 3 The RV variability of giants and their Li content
Figure 5: The fraction of RV-variable stars in our samples of Li-normal and Li-rich giants. Insofar as the high-variability classes 5a-5b are indicative of binaries, the fraction of giants in the combined 5a-5b class can be thought of as the close binary fraction for the corresponding sample. Left: No distinction is made between the evolutionary phase of the giants. Center: only first-ascent RGB giants. Right: only RC giants. Evolutionary phases were determined following Martell et al. (2021). There is a slight hint among first ascent (RGB) giants of higher RV variability (higher close binary fraction) for Li-rich than for Li-normal giants, although marginal given the estimated uncertainties, When looking at the RC giants alone, however, the fraction of RV variables is the same among Li-rich and Li-normal stars.
@@ 3.1 RV variability and Li abundance
We applied our classification scheme of Section 2.1 both to the 5667 giants with three RV epochs, and to the subsample of 1418 with good Li determinations from GALAH DR3. We used $RV_{1}=RV_{\mathrm{Gaia}}$, $RV_{2}=RV_{\mathrm{GALAH}}$, and $RV_{3}=RV_{\mathrm{RAVE}}$. Table 2 shows the fraction of giants in each variability class for both samples. We noted that the proportion of stars in each class is very similar for both subsets, i.e., the giants with A(Li) measurement are a representative sample of the total set of giants with RV measurements registered in Gaia, GALAH, and RAVE. In fact, even when the errors of the values in Table 2 were taken to be a Poisson uncertainty of the counts for simplicity, all the fractions of the classes for the giants with A(Li) are within $<1\sigma$ of the fractions in the sample with all the giants.
An important caveat to consider is that some giants can exhibit pulsations, which can produce variations in radial velocities up to $\lesssim 1\ \mathrm{km\ s^{-1}}$ and with typical values ranging from $\sim 10$ to $100\ \mathrm{m\ s^{-1}}$ (e.g. Hekker et al. 2008). We note that $\Delta{RV_{\mathrm{max}}}=359\ \mathrm{m\ s^{-1}}$ is the minimum variation in our giants in classes 5a-5b, with only about $1.1\%$ and $8.4\%$ of them showing variations of $\Delta{RV_{\mathrm{max}}}<500\ \mathrm{m\ s^{-1}}$ and $\Delta{RV_{\mathrm{max}}}<1\ \mathrm{km\ s^{-1}}$, respectively. Thus, while we cannot conclusively distinguish between RV variations caused by pulsations and those induced by companion stars, only a small fraction are likely related to pulsations among our highly variable giants. A combination of both processes can also contribute to RV variations (Beck et al. 2014). Moreover, Cepheids and RRLyrae objects are among the most variable stars due to radial pulsations, with amplitudes $>10\ \mathrm{km\ s^{-1}}$. Given that Gaia DR3 provides a catalog of such objects (Ripepi et al. 2023), we verified that our 1418 giants with measured Li abundances are not part of this catalog.
Table 2: Percentage of giants in each variability class for the sample with all the giants with RV measurements in Gaia, GALAH, and RAVE (5667 giants), and for the subsample with A(Li) measurements in GALAH (1418 giants).
@@
Now, we combine the A(Li) information with the variability classes obtained to explore if there is any relation between Li content and RV variability. We illustrate in Figure 5 the fractions of giants as a function of our variability classes, for Li-normal (A(Li)$\leq$1.5) and Li-rich giants separately [^2] We included the level of uncertainty of our variability classification by considering the results of the tests made in Section 2.2. Since we found about $p_{nv}=0.802$ and $p_{v}=0.849$ of effectiveness for classifying non-variables and variables respectively, we calculate the expected value of correctly detected non-variables to be $p_{nv}n_{nv}$ for $n_{nv}$ giants found in classes $\leq 4$. Consequently, $p_{v}n_{v}$ is the value of correctly detected variables for $n_{v}$ giants found in classes 5a-5b. These values of effectiveness have an intrinsic uncertainty associated. We add that contribution to the error of the expected values for non-variables or variables by taking $\sigma_{g}=\sqrt{\sigma_{e}^{2}+\sigma_{p}^{2}}$, where $\sigma_{g}$ is the error on the expected number of giants in a certain variability type, $\sigma_{e}=\sqrt{pn}$ the Poisson error for $n$ measurements with individual probabilities $p$ to obtain the expected value, and $\sigma_{p}=ne_{p}$ for $e_{p}$ the error of $p$. From the tests presented in Section 2.2 we estimated $e_{p,nv}=0.05$ for the non-variables and $e_{p,v}=0.16$ for the variables.. In this case, to avoid clutter, we consider joint classes, i.e., class 5a-5b, class 4-3, and class 2-1 (for simplicity we usually refer in this analysis as non-variables the giants in classes $\leq 4$, and variables to those giants with high certainty of variability, i.e., classes 5a-5b). In the left-hand panel, we present the expected fraction of giants in each variability class for the Li-normal and Li-rich giants without considering the evolutionary phase. It can be seen that the fraction of Li-rich giants is essentially the same as that of Li-normal giants for all RV variability classes, given the error bars.
@@
In the center panel of Figure 5 is the expected fraction of RGB stars in each variability class. In this case, the Li-rich RGB population has a larger fraction of stars in the high variability class than the $\sim 33\%$ found in the Li-normal RGB and the samples without considering the evolutionary phase. Consequently, the fraction of non-variable Li-rich giants is smaller than those in the other Li-normal samples. However, the total number of Li-rich RGBs is the smallest sample analyzed in Figure 5, and as such the errors of the fractions are also large. This means that the error bars of the fractions in each variability class are blended at least within one standard deviation with the non-Li-rich RGBs. This apparent difference, however, goes away completely among the subsample of RC giants (right-hand panel), and we note that not only is the fraction of Li-rich RC giants the same as that of Li-normal RC giants for all RV-variability classes, but the fraction of Li-rich giants with large RV variability has decreased from 52% in the RGB to 30% in the RC. Even though our estimated error bars may account for this difference between Li-rich giants in the RGB and the RC, we think it is something worth following up with further work.
@@
Figure 6: Left Panel: A(Li) distribution for each Class of variability obtained for the giants. $\mathrm{N_{s}/N_{class}}$ is the fraction of stars in a class for the corresponding bin of A(Li). The percentages represent the fraction in each class just like in Table 2. Right Panel: A(Li) distribution for each joint Class of variability obtained for the Li-rich (A(Li)¿1.5) giants. For this case, the variability is high for class 5a-5b, medium for class 4-3, and low for class 2-1; the percentages represent the fraction of giants in each joint class for the total Li-rich giants with errors obtained following the same method applied in Table 2.
@@
To get greater insight into the relationship between A(Li) in giants and variability we analyze the A(Li) distribution for all the variability classes defined under the $w$-criteria. The left panel of Figure 6 shows the A(Li) distribution of each class obtained from the $w_{12}$-$w_{t}$ criterion for our sample of giants. The vertical axis shows the number of stars $\mathrm{N_{s}}$ in a given A(Li) bin of range 0.5 dex (starting from A(Li)=$-1$), normalized by $\mathrm{N_{class}}$, the total number of stars with Li determination that are in the corresponding variability class. The first immediate conclusion from Figure 6 is that all variability classes follow almost the same distribution, which is very similar to a normal distribution. Moreover, they all reach their maximum value at the same bin of Li abundance: A(Li) centered at 0.75 dex. This means that there are no obvious differences regarding Li content among all the variability classes, which we are using as a proxy for the likelihood of binarity. There is a hint that the peak of the A(Li) distribution may be slightly wider for the most variable classes (4, 5a, and 5b). But, given the size of the error bars, these differences are not larger than $1\sigma$, and therefore not significant.
@@
The right-hand panel of Figure 6 zooms into the high-end tail of the A(Li) distributions from the left-hand panel, starting from A(Li)=1.5. In this panel, we consider joint classes. Thus, $\mathrm{N_{s}/N_{class}}$ is the fraction of stars in a joint class for a given range of A(Li). Here we consider the fractions in the bins normalized with respect to the Li-rich giants in each class. Consequently, given the different normalizations, the shape of the distributions exhibits slight variations compared to what is observed in the left-hand panel of the figure. Giants in class 5a-5b are therefore those Li-rich giants with high certainty of being in a binary system (because they have the largest variation in RV), and they correspond to $37.6\%$ of all the Li-rich stars in our sample. This is only about $5\%$ larger than the fraction of stars with high certainty of variability when considering all giants with A(Li) measurement ($32.9\%$) and all giants with RV in the three surveys ($32.3\%$). Considering error bars, $33\%$ is about one standard deviation from the fraction of Li-rich giants in class 5a-5b (37.6%) which is marginal. Therefore, regardless of their Li abundance, there are about one-third of giants with a high probability of being in a binary system. At the same time, among the Li-rich giants only, there are no significant differences among the A(Li) distributions, they all show that the fraction $\mathrm{N_{s}/N_{class}}$ of giants in each A(Li) bin is almost the same for all the variability classes.
@@
These results are consistent with previous work with smaller samples of red giants but more RV measurements per star. Jorissen et al. (2020) followed up a set of 24 red giants for about 9 years and less than $36\%$ of the Li-rich giants were identified with a high probability of being in binary systems under the F2-distribution method. Aguilera-Gómez et al. (2023) also compared the A(Li) distribution of a handful of field red giants, of which some were confirmed as spectroscopic binaries, and others with a planet, and they also did not find any trend between a companion and the Li enrichment. Other correlations have been studied that could hint towards binary interaction producing very Li-rich giants, e.g., with [[[[C/N]]]] ratio (Tayar et al. 2023). This would be an additional method that could be explored to understand if binary interactions are responsible for all or a subset of Li-rich giants. Additionally, extensive follow-up of Li-rich giants and measurements of multiple RV epochs is necessary to confirm the results presented in this work.
@@
While A(Li)¿1.5 is the typical threshold for considering Li-rich giants, there is a fraction of objects that have Li abundances higher than the expected value from the Big Bang nucleosynthesis, i.e., A(Li)¿2.7 (Cyburt et al. 2008; Fields et al. 2020), and some with higher values than the Solar System meteoritic value (Lodders & Fegley 1998), i.e., A(Li)¿3.3. These are often referred to as super Li-rich giants. Such abundances could require internal Li production to be reached (or probably a combination of internal and external processes). The interaction with a binary companion has been proposed as one of the processes that can favor the Li production and transport to the surface in giants (Casey et al. 2019), then it is worth studying the fraction of super Li-rich giants in our sample.
Table 3: Number of giants identified as Class 5a-5b and Class $\leq$ 4 for different Li-rich ranges. The percentages in parenthesis show the fraction of giants in the class with respect to the total number of giants in the range of A(Li). The table was split into three sections to also identify the fraction of Li-rich giants in each class of variability in a specific evolutionary phase (RGB or RC) using the crossmatch of the classified GALAH DR3 giants of Martell et al. (2021).
@@
Table 3 shows the amount of Li-rich and super Li-rich giants that are identified as Class 5a-b and Class $\leq 4$ for the total sample and those classified as RGB and RC separately. We found that the fraction of giants with high variability among super Li-rich giants with A(Li)¿2.7 is also about one-third, and it is ¡$25\%$ for those with A(Li)¿3.3 without taking into account the evolutionary phase, which is consistent with what we discussed in Section 3.1 where independent of the range of Li considered the fraction of giants with high variability is in a tight range.
@@
These results require further investigation since the total number of super Li-rich is small (especially for the RGB case). Note, for example if we consider Poisson deviations in the data of Table 3, the errors of the fractions of super Li-rich RGB (A(Li)¿2.7 and ¿3.3) are about $15\%$-$30\%$. Further, we previously estimated that about 9 of the Li-rich giants in our sample could be in the AGB phase. Some of them are likely to be super Li-rich since depending on the mass, the extension of the convective envelope during the AGB pulsating phase can rapidly transport the Li to the surface, allowing it to reach values A(Li)$>2.7$ (see Ventura & D’Antona 2010; Lau et al. 2012).
@@ 4 Binarity and Variability in Gaia DR3
In this section, we explore the feasibility of our methods for detecting RV variability and the results when analyzing the A(Li) distribution of different classes using indicators of binarity and RV variation reported in the last data release of Gaia.
@@
Gaia DR3 offers an identification for non-single systems primarily based on orbital resolution (Hambly et al. 2022). A non_single_star>0 (NSS¿0) flag in Gaia indicates that the methods have detected a binary system. We tested this flag in our set of 1418 and obtained that $5.2\%$ of them have NSS¿0, significantly different from the $32.9\%$ with high variability based on our $w$ values. Gaia Collaboration et al. (2022) and Pourbaix et al. (2022) have pointed out that the NSS catalog of Gaia is still incomplete, so we need to keep this caveat into account. We estimate a degree of incompleteness by testing the NSS flag on our SB9 stars and found that only $60.6\%$ of them have NSS¿0. Then, based on the tests made in Section 2, our $w$ parameters are likely classifying more close binaries than the NSS flag in Gaia DR3. In Panel (i) of Figure 7 we also show the Li abundance distributions for single stars and binary systems found with the NSS flag. The two distributions are alike, except that almost all the Li-rich giants are not binaries and there are no NSS¿0 systems with A(Li)¿1.8. Even considering the incompleteness of the NSS flag, these results suggest an agreement with what we obtained using the $w$ parameters, i.e., Li-rich giants should not be mostly related to binary systems.
@@
An alternative way of detecting RV variability from parameters in the Gaia DR3 catalog was reported by Katz et al. (2022): a star with normalized goodness of fit (GoF) rv_renormalised_gof > 4, and consistency for RV measurements rv_chisq_pvalue $\leq$ 0.01 ($\leq$0.01), is variable [^3] The GoF describes an empirical quantity that represents the scatter of RVs when compared to the typical epoch uncertainty. p is the probability that the $\chi^{2}$ for the RVs exceeds an expected value of the probability distribution with certain significance (see Hambly et al. 2022).. When testing these criteria in our SB9 stars, $92.0\%$ of them are identified as variables. This is a similar performance as the $w$-criteria for detecting binaries ($\sim 85$-$94\%$). We use this alternate method for our giants and compare the A(Li) distributions obtained for variables and non-variables in the second panel (ii) of Figure 7. These distributions appear to be quite similar to what is shown in panel (i). Furthermore, just $9.8\%$ of the giants were identified as variables, and the maximum Li abundance identified for this subsample is A(Li)$\sim$ 2.7. This disagrees with the about $33\%$ of variable giants found by our method and also by Badenes et al. (2018) when analyzing APOGEE giants.

### v3

@@ 3 The RV variability of giants and their Li content
Figure 5: Fraction of RV-variable stars in our samples of Li-normal and Li-rich giants. Insofar as the high-variability classes 5a-5b are indicative of binaries, the fraction of giants in the combined 5a-5b class can be thought of as the close binary fraction for the corresponding sample. In the left panel, no distinction is made between the evolutionary phases of the giants. The center panel shows only the first-ascent RGB giants, while the right panel displays only the RC giants. Evolutionary phases were determined following Martell et al. (2021). There is a slight hint among first ascent (RGB) giants of higher RV variability (higher close binary fraction) for Li-rich than for Li-normal giants, although marginal given the estimated uncertainties, When looking at the RC giants alone, however, the fraction of RV variables is the same among Li-rich and Li-normal stars.
@@ 3.1 RV variability and Li abundance
We applied our classification scheme of Sect. 2.1 both to the 5667 giants with three RV epochs, and to the subsample of 1418 with good Li determinations from GALAH DR3. We used $RV_{1}=RV_{\mathrm{\textit{Gaia}}}$, $RV_{2}=RV_{\mathrm{GALAH}}$, and $RV_{3}=RV_{\mathrm{RAVE}}$. Table 2 shows the fraction of giants in each variability class for both samples. We noted that the proportion of stars in each class is very similar for both subsets, that is, the giants with A(Li) measurement are a representative sample of the total set of giants with RV measurements registered in Gaia, GALAH, and RAVE. In fact, even when the errors of the values in Table 2 were taken to be a Poisson uncertainty of the counts for simplicity, all the fractions of the classes for the giants with A(Li) are within $<1\sigma$ of the fractions in the sample with all the giants.
An important caveat to consider is that some giants can exhibit pulsations, which can produce variations in RVs up to $\lesssim 1\ \mathrm{km\ s^{-1}}$ and with typical values ranging from $\sim 10$ to $100\ \mathrm{m\ s^{-1}}$ (e.g., Hekker et al. 2008). We note that $\Delta{RV_{\mathrm{max}}}=359\ \mathrm{m\ s^{-1}}$ is the minimum variation in our giants in classes 5a-5b, with only about $1.1\%$ and $8.4\%$ of them showing variations of $\Delta{RV_{\mathrm{max}}}<500\ \mathrm{m\ s^{-1}}$ and $\Delta{RV_{\mathrm{max}}}<1\ \mathrm{km\ s^{-1}}$, respectively. Thus, while we cannot conclusively distinguish between RV variations caused by pulsations and those induced by companion stars, only a small fraction are likely related to pulsations among our highly variable giants. A combination of both processes can also contribute to RV variations (Beck et al. 2014). Moreover, Cepheids and RRLyrae objects are among the most variable stars due to radial pulsations, with amplitudes $>10\ \mathrm{km\ s^{-1}}$. Given that Gaia DR3 provides a catalog of such objects (Ripepi et al. 2023), we verified that our 1418 giants with measured Li abundances are not part of this catalog.
Table 2: Percentage of giants in each variability class.
@@
 [^3] The sample includes all the giants with RV measurements in Gaia, GALAH, and RAVE (5667 giants), and for the subsample with A(Li) measurements in GALAH (1418 giants).
@@
Now, we combine the A(Li) information with the variability classes obtained to explore if there is any relation between Li content and RV variability. We illustrate in Fig. 5 the fractions of giants as a function of our variability classes, for Li-normal (A(Li)$\leq$1.5) and Li-rich giants separately [^4] We included the level of uncertainty of our variability classification by considering the results of the tests made in Sect. 2.2. Since we found about $p_{nv}=0.802$ and $p_{v}=0.849$ of effectiveness for classifying non-variables and variables respectively, we calculate the expected value of correctly detected non-variables to be $p_{nv}n_{nv}$ for $n_{nv}$ giants found in classes $\leq 4$. Consequently, $p_{v}n_{v}$ is the value of correctly detected variables for $n_{v}$ giants found in classes 5a-5b. These values of effectiveness have an intrinsic uncertainty associated. We add that contribution to the error of the expected values for non-variables or variables by taking $\sigma_{g}=\sqrt{\sigma_{e}^{2}+\sigma_{p}^{2}}$, where $\sigma_{g}$ is the error on the expected number of giants in a certain variability type, $\sigma_{e}=\sqrt{pn}$ the Poisson error for $n$ measurements with individual probabilities $p$ to obtain the expected value, and $\sigma_{p}=ne_{p}$ for $e_{p}$ the error of $p$. From the tests presented in Sect. 2.2 we estimated $e_{p,nv}=0.05$ for the non-variables and $e_{p,v}=0.16$ for the variables.. In this case, to avoid clutter, we consider joint classes, that is, class 5a-5b, class 4-3, and class 2-1 (for simplicity we usually refer in this analysis as non-variables the giants in classes $\leq 4$, and variables to those giants with high certainty of variability, i.e., classes 5a-5b). In the left-hand panel, we present the expected fraction of giants in each variability class for the Li-normal and Li-rich giants without considering the evolutionary phase. It can be seen that the fraction of Li-rich giants is essentially the same as that of Li-normal giants for all RV variability classes, given the error bars.
@@
In the center panel of Fig. 5 is the expected fraction of RGB stars in each variability class. In this case, the Li-rich RGB population has a larger fraction of stars in the high variability class than the $\sim 33\%$ found in the Li-normal RGB and the samples without considering the evolutionary phase. Consequently, the fraction of non-variable Li-rich giants is smaller than those in the other Li-normal samples. However, the total number of Li-rich RGBs is the smallest sample analyzed in Fig. 5, and as such the errors of the fractions are also large. This means that the error bars of the fractions in each variability class are blended at least within one standard deviation with the non-Li-rich RGBs. This apparent difference, however, goes away completely among the subsample of RC giants (right-hand panel), and we note that not only is the fraction of Li-rich RC giants the same as that of Li-normal RC giants for all RV-variability classes, but the fraction of Li-rich giants with large RV variability has decreased from 52% in the RGB to 30% in the RC. Even though our estimated error bars may account for this difference between Li-rich giants in the RGB and the RC, we think it is something worth following up with further work.
@@
Figure 6: The left panel shows the A(Li) distribution for each class of variability obtained for the giants. We note that $\mathrm{N_{s}/N_{class}}$ is the fraction of stars in a class for the corresponding bin of A(Li). The percentages represent the fraction in each class just like in Table 2. The right panel shows the A(Li) distribution for each joint class of variability obtained for the Li-rich (A(Li)¿1.5) giants. For this case, the variability is high for class 5a-5b, medium for class 4-3, and low for class 2-1; the percentages represent the fraction of giants in each joint class for the total Li-rich giants with errors obtained following the same method applied in Table 2.
@@
To get greater insight into the relationship between A(Li) in giants and variability we analyze the A(Li) distribution for all the variability classes defined under the $w$-criteria. The left panel of Fig. 6 shows the A(Li) distribution of each class obtained from the $w_{12}$-$w_{t}$ criterion for our sample of giants. The vertical axis shows the number of stars $\mathrm{N_{s}}$ in a given A(Li) bin of range 0.5 dex (starting from A(Li)=$-1$), normalized by $\mathrm{N_{class}}$, the total number of stars with Li determination that are in the corresponding variability class. The first immediate conclusion from Fig. 6 is that all variability classes follow almost the same distribution, which is very similar to a normal distribution. Moreover, they all reach their maximum value at the same bin of Li abundance: A(Li) centered at 0.75 dex. This means that there are no obvious differences regarding Li content among all the variability classes, which we are using as a proxy for the likelihood of binarity. There is a hint that the peak of the A(Li) distribution may be slightly wider for the most variable classes (4, 5a, and 5b). But, given the size of the error bars, these differences are not larger than $1\sigma$, and therefore not significant.
@@
The right-hand panel of Fig. 6 zooms into the high-end tail of the A(Li) distributions from the left-hand panel, starting from A(Li)=1.5. In this panel, we consider joint classes. Thus, $\mathrm{N_{s}/N_{class}}$ is the fraction of stars in a joint class for a given range of A(Li). Here we consider the fractions in the bins normalized with respect to the Li-rich giants in each class. Consequently, given the different normalizations, the shape of the distributions exhibits slight variations compared to what is observed in the left-hand panel of the figure. Giants in class 5a-5b are therefore those Li-rich giants with high certainty of being in a binary system (because they have the largest variation in RV), and they correspond to $37.6\%$ of all the Li-rich stars in our sample. This is only about $5\%$ larger than the fraction of stars with high certainty of variability when considering all giants with A(Li) measurement ($32.9\%$) and all giants with RV in the three surveys ($32.3\%$). Considering error bars, $33\%$ is about one standard deviation from the fraction of Li-rich giants in class 5a-5b (37.6%) which is marginal. Therefore, regardless of their Li abundance, there are about one-third of giants with a high probability of being in a binary system. At the same time, among the Li-rich giants only, there are no significant differences among the A(Li) distributions, they all show that the fraction $\mathrm{N_{s}/N_{class}}$ of giants in each A(Li) bin is almost the same for all the variability classes.
@@
These results are consistent with previous work with smaller samples of red giants but more RV measurements per star. Jorissen et al. (2020) followed up a set of 24 red giants for about 9 years and less than $36\%$ of the Li-rich giants were identified with a high probability of being in binary systems under the F2-distribution method. Aguilera-Gómez et al. (2023) also compared the A(Li) distribution of a handful of field red giants, of which some were confirmed as spectroscopic binaries, and others with a planet, and they also did not find any trend between a companion and the Li enrichment. Other correlations have been studied that could hint toward binary interaction producing very Li-rich giants, for example, with [[[[C/N]]]] ratio (Tayar et al. 2023). This would be an additional method that could be explored to understand if binary interactions are responsible for all or a subset of Li-rich giants. Additionally, extensive follow-up of Li-rich giants and measurements of multiple RV epochs is necessary to confirm the results presented in this work.
We also tested our criteria for RV variability using the 24 giants from Jorissen et al. (2020). In their work, each star has multiple RV measurements. For each giant, we designated the measurement with the largest RV error as $RV_{3}$ and explored all possible combinations with the remaining RVs to determine $RV_{1}$ and $RV_{2}$. Among the 11 definitive and possible binaries identified in their sample, 9 were consistently classified as class 5a-5b in more than 50% of their RV combinations, with 7 classified as such in over 90% of combinations. For the 13 non-variables, all were classified in classes $\leq$4 in more than 50% of their RV combinations, with 8 classified as such in over 90% of combinations. This result indicates that our variability criteria can reproduce similar results to previous studies, although limited by the number of measurements used for our classifications.
@@
While A(Li)¿1.5 is the typical threshold for considering Li-rich giants, there is a fraction of objects that have Li abundances higher than the expected value from the Big Bang nucleosynthesis, that is, A(Li)¿2.7 (Cyburt et al. 2008; Fields et al. 2020), and some with higher values than the Solar System meteoritic value (Lodders & Fegley 1998), that is, A(Li)¿3.3. These are often referred to as super Li-rich giants. Such abundances could require internal Li production to be reached (or probably a combination of internal and external processes). The interaction with a binary companion has been proposed as one of the processes that can favor the Li production and transport to the surface in giants (Casey et al. 2019), then it is worth studying the fraction of super Li-rich giants in our sample.
Table 3: Giants in class 5a-5b and class $\leq$ 4 for different Li-rich ranges.
@@
 [^5] The number of giants in each class is presented, while the percentages in parentheses indicate the fraction of giants in each class relative to the total number of giants in the range of A(Li). The table is divided into three sections to identify the number of Li-rich giants in each variability class during specific evolutionary phases (RGB or RC), based on the cross-match with the classified GALAH DR3 giants from Martell et al. (2021).
Table 3 shows the amount of Li-rich and super Li-rich giants that are identified as class 5a-b and class $\leq 4$ for the total sample and those classified as RGB and RC separately. We found that the fraction of giants with high variability among super Li-rich giants with A(Li)¿2.7 is also about one-third, and it is ¡$25\%$ for those with A(Li)¿3.3 without taking into account the evolutionary phase, which is consistent with what we discussed in Sect. 3.1 where independent of the range of Li considered the fraction of giants with high variability is in a tight range.
@@
These results require further investigation since the total number of super Li-rich giants is small (especially for the RGB case). It is important to note, for example, that if we consider Poisson deviations in the data of Table 3, the errors of the fractions of super Li-rich RGB (A(Li)¿2.7 and ¿3.3) are about $15\%$-$30\%$. Further, we previously estimated that about 9 of the Li-rich giants in our sample could be in the AGB phase. Some of them are likely to be super Li-rich since depending on the mass, the extension of the convective envelope during the AGB pulsating phase can rapidly transport the Li to the surface, allowing it to reach values A(Li)$>2.7$ (see Ventura & D’Antona 2010; Lau et al. 2012).
@@ 4 Binarity and variability in Gaia DR3
In this section we explore the feasibility of our methods for detecting RV variability and the results when analyzing the A(Li) distribution of different classes using indicators of binarity and RV variation reported in the last data release of Gaia.
@@
Gaia DR3 offers an identification for non-single systems primarily based on orbital resolution (Hambly et al. 2022). A non_single_star>0 (NSS¿0) flag in Gaia indicates that the methods have detected a binary system. We tested this flag in our set of 1418 and obtained that $5.2\%$ of them have NSS¿0, significantly different from the $32.9\%$ with high variability based on our $w$ values. Gaia Collaboration et al. (2023) and Pourbaix et al. (2022) have pointed out that the NSS catalog of Gaia is still incomplete, so we need to keep this caveat into account. We estimate a degree of incompleteness by testing the NSS flag on our SB9 stars and found that only $60.6\%$ of them have NSS¿0. Then, based on the tests made in Sect. 2, our $w$ parameters are likely classifying more close binaries than the NSS flag in Gaia DR3. In Panel (i) of Fig. 7 we also show the Li abundance distributions for single stars and binary systems found with the NSS flag. The two distributions are alike, except that almost all the Li-rich giants are not binaries and there are no NSS¿0 systems with A(Li)¿1.8. Even considering the incompleteness of the NSS flag, these results suggest an agreement with what we obtained using the $w$ parameters, that is, Li-rich giants should not be mostly related to binary systems.
@@
An alternative way of detecting RV variability from parameters in the Gaia DR3 catalog was reported by Katz et al. (2023): a star with normalized goodness of fit (GoF) rv_renormalised_gof > 4, and consistency for RV measurements rv_chisq_pvalue $\leq$ 0.01 ($\leq$0.01), is variable [^6] The GoF describes an empirical quantity that represents the scatter of RVs when compared to the typical epoch uncertainty. We note that p is the probability that the $\chi^{2}$ for the RVs exceeds an expected value of the probability distribution with a certain significance (see Hambly et al. 2022).. When testing these criteria in our SB9 stars, $92.0\%$ of them are identified as variables. This is a similar performance as the $w$-criteria for detecting binaries ($\sim 85$-$94\%$). We use this alternate method for our giants and compare the A(Li) distributions obtained for variables and non-variables in the second panel (ii) of Fig. 7. These distributions appear to be quite similar to what is shown in panel (i). Furthermore, just $9.8\%$ of the giants were identified as variables, and the maximum Li abundance identified for this subsample is A(Li)$\sim$ 2.7. This disagrees with the about $33\%$ of variable giants found by our method and also by Badenes et al. (2018) when analyzing APOGEE giants.
The discrepancies between our w-criteria and the method of Katz et al. (2023) are likely due to the tight criteria for variability the latter considers: it requires analyzing more than ten epochs. Thus, to relax these criteria to something more comparable to our $w$ method, we obtained the values of p and GoF using the 3 epochs of RV used in Sect. 3 and following the description of Hambly et al. (2022). In doing so, similarly to the previous tests, 10.4% of the giants were identified as variables. Furthermore, the formulation of the GoF is almost the same as the F2 distribution (e.g., see Jofré et al. 2023) [^7] Except that we should include a normalization factor mode(UWE)$\sim$ 1 for a range $\mathrm{T_{eff}}\approx 4000$-$5800$ K (see Sartoretti et al. 2022).. Then, by relaxing p in the description of Katz et al. (2023) we should obtain a proxy of the F2 distribution. We noted that as we relax p, we classify more giants as variables, and when using just the GoF, we obtain 38.2% of the giants being variables and with an A(Li) distribution almost identical to what we obtained when using the w-criteria.